OpenStudy (idkwut):

The decibel level of a sound is given by the equation D = 10log10 (I/I_0) where I is the sound's intensity and I0 corresponds to the intensity of the weakest sound that can be detected by the human ear. A sound of 140 decibels produces pain in the average human ear. Approximately how many times greater than I0 must the intensity I of a sound be to reach this decibel level?

OpenStudy (idkwut):

a. 22026 times greater b. 14 times greater c. 1.14612 times greater d. 2.63906 times greater e. 1014 times greater

OpenStudy (idkwut):

@zepdrix @Compassionate Help plz?!

OpenStudy (idkwut):

Opps, I think it ment I_0, not 10

zepdrix (zepdrix):

So our function, decibels as a function of sound intensity,\[\Large\bf\sf \mathcal D\quad=\quad 10\log\left(\frac{\mathcal I}{\mathcal I_o}\right)\]

zepdrix (zepdrix):

And they want us to compare I to I_o when D=140

zepdrix (zepdrix):

\[\Large\bf\sf 140\quad=\quad 10\log\left(\frac{\mathcal I}{\mathcal I_o}\right)\]

zepdrix (zepdrix):

So we want to start by solving for \(\Large\bf\sf \dfrac{\mathcal I}{\mathcal I_o}\) How can we do that? :)

OpenStudy (idkwut):

Do we multiply both sides

zepdrix (zepdrix):

Umm let's divide each side by 10, that will clean things up a little bit.

zepdrix (zepdrix):

\[\Large\bf\sf 14\quad=\quad \log\left(\frac{\mathcal I}{\mathcal I_o}\right)\]

zepdrix (zepdrix):

From here you can either `exponentiate` or simply rewrite it in exponential form if you remember how.

OpenStudy (idkwut):

Oh sry, I meant to put 10^14 not 1014 on the choice e.

OpenStudy (idkwut):

I don't know how to do tht

zepdrix (zepdrix):

So we'll try exponentiating. It's a bit of a weird process but worth getting comfortable with. Rewrite each side as an exponent of base 10,\[\huge\bf\sf 10^{14}\quad=\quad 10^{\log\left(\frac{\mathcal I}{\mathcal I_o}\right)}\]

zepdrix (zepdrix):

Since log base 10, and the exponential function base 10 are `inverse functions` of one another. They "undo" each other. You can think of it as "cancelling out" if that's easier to remember. It's not as accurate, but that's ok.\[\Large\bf\sf 10^{14}\quad=\quad \left(\frac{\mathcal I}{\mathcal I_o}\right)\]

zepdrix (zepdrix):

Then multiply I_o to the other side,

zepdrix (zepdrix):

\[\Large\bf\sf 10^{14}\mathcal I_o\quad=\quad \mathcal I\]